Sol Summary – February 13th

Sol: 3

Summary Title: The sound of Martian winds

Author’s name: Natalia Larrea

Mission Status: Nominal

Sol Activity Summary: We woke up with the smell of home-made brownies and coffee. The first EVA left at 10.00 am driving Curiosity and Opportunity rovers heading to Candor Chasma. The team collected excellent geological samples and had the chance to explore the area. Meanwhile, at the Hab, the team worked on new maintenance tasks. We checked the status of the EVA suits making sure they are fully operational. Batteries were replaced in those suits not working properly, making them now ready to be used. At the return of the EVA team, all the crew enjoyed together some Martian pasta with dried tomato and chicken sauce. In the afternoon, the second EVA team headed to explore the area surrounding the Hab.

The plan was to test Crew 205’s drone, collect samples and conduct photography. However, the Martian weather complicated the task. Winds were too strong to fly the drone. The team continued nonetheless, with its plans of climbing the Old Repeater Point. The view of MDRS was great from there! However, in the field, the rain started to hit and the space helmets started to become foggy. Walking on the hilly surface in such conditions was not easy. The team then decided to head back to the Hab, enjoy a warm coffee and tea, and continue working there. Tonight we plan movie night … maybe “The Martian”?;))

Look Ahead Plan: Most of the day tomorrow will be dedicated to working “from home” on the different projects. After a few days on Mars, we have started collecting many pieces of great science and data which need to be analyzed. A team will also head the Old Repeater Point and North Ridge for a short test the drone (if the weather permits), continuing with the initial plan of today.

Tomorrow is also Saint Valentine’s day so we plan a special evening and dinner with “Secret Valentine” presents included. In addition, it will be the first day we start rotating roles in the crew. The members in the position of commander and XO will be rotating for the next week.

Anomalies in work: None

Weather: Cloudy all day long, windy with light rain in the afternoon

Crew Physical Status: Nominal

EVA#05: 0520300/4241100 (Candor Chasma)
EVA#06: 4250900/517950 (Old Repeater Point), 4251500/518600 (North Ridge, not reached)

Reports to be filed: Sol Summary, EVA report, EVA request, Operations Report, HSO Report, Astronomy Report, Science Report, Greenhab Report, Journalist Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site